OpenStudy (anonymous):

how to solve x/8+25=0

OpenStudy (anonymous):

subtract 25 then multiply by 8

OpenStudy (anonymous):

that's undefined x=0

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

x/8+25=0 x/8+25-25=0-25 x/8=-25 x/8*8=-25*8 x=-200

OpenStudy (anonymous):

Is the equation like this? \[\frac{ x }{ 8 }+25=0\] or \[\frac{ x }{ 8+25 }=0\]

OpenStudy (anonymous):

For the first one, x=-200.

OpenStudy (anonymous):

yes that is correct i agree with @AuroraShay

OpenStudy (anonymous):

@AuroraShay could you help me with my question please

OpenStudy (anonymous):

Thank you mathman